I've continued to work on my Gaussianity tests for S5 L1 data.
Following the statistical measure in Ando et al. (2003), I've calculated the Laguerre coefficient, c2, for all frequencies present in my S5 L1 PSD as a metric of Gaussianity. When c2 is zero, the distribution is Gaussian. A positive c2 corresponds to glitchy noise, while a negative c2 suggests stationary noise.
Below is a plot displaying variation in c2 for this PSD:
By observing the c2 value and histogram of distribution of various PSD values at a given frequency, we can elucidate statistical differences in the Gaussian nature of noise at that frequency which are unclear in the standard PSD.
In my last post I calculated the different subtractions (offline) that could be done to YARM alone just to get a sense of what seismometers were better witnesses for the Wiener filter calculation.
In this eLOG I show what subtractions can be done when the MCL has FF on (as well as Eric's PRC FF), with the SISO filter described on elog:11496.
The plot below shows what can be done offline,
What is great about this results is that the T240-X and T240-Y channels are plenty enough to mitigate any remaining YARM seismic noise but also to get rid of that nasty peak at 55 Hz induced by the MCL FF filter.
The caviat, I haven't measured the TF for the ETMY actuator to YARM control signal. I need to do this and recompute the FIR filters with the prefiltered witnesses in order to move on to the IIR converions and online FF!
I fixed a shaker that was claimed to be broken. I had to cut the rubber membrane to open the head.
Once it was opened, the cause of the trouble was obvious. The soldering joint could not put up with the motion of the head.
It is interesting to see that the spring has the damping layer between the metal sheets.
After the repair the DC resistance was measured. It was 1.9Ohm. The side of the shaker chassis said "3.5Ohm, Max 15VA". So it can take more than 4A (wow).
I gave 2A DC from the bench top supply and turn the current on and off. I could confirm the head was moving.
I'll claim the use of this shaker for the seismometer development.
Plotte below are the resultant subtractions for YARM using different witness configurations,
The best subtraction happens with all the channels of both the GUR1 and T240 seismometers, but one gets just as good subtraction without using the z channels as witnesses.
Also, why is the T240 seismometer better at subtracting noise for YARM compared to what GUR1 alone can acomplish? Using only the X and Y channels for the T240 gave the third best subtraction(purple trace).
My plan for now is as follows:
1) Measure the transfer function from the ETMY actuator to the YARM control signal
2) Collect data for YARM when FF for MCL is on in order to see what kind of subtractions can be done.
IIRC, the Guralp box's 3rd set of channels do not have all of the modifications that were made on channels 1 and 2.
Let's dismantle the I/F unit from the rack and connect the cable with the lid open.
We need to trace the signal.
Atm1, Before cable swap
Atm2, The long cables were swapped at the input of the interface box.
We can conclude that the problem is in the interface box
I wonder if interface box input 3 is wired?
I took data from 1123495750 to 1123498750 GPS time (Aug 13 at 3AM, 50 mins of data) for C1:LSC-YARM_OUT_DQ, and all T240 and GUR1 channels.
Here is the PSD of the YARM_OUT, showing the data that I will use to train the FIR filter:
Coherence plots for YARM and all channels of T240 and GUR1 sesimometers are shown below. This will help determine what regions to preweight the best before computing FIR filter. They also show how GUR1 is back to work compared to those of elog:11457.
The mode cleaner FF static filtering is by no means done. More work has to be done in order to succefuly implement it, by the means of fine tuning the IIR fit and finding better MISO Wiener filters.
I have begun to look at implementing FF to the YARM cavity for several reasons.
1) Even if the mode cleaner FF is set up as best as we can, there will still be seismic noise coupling into the arm cavities.
2) YARM is in the way of the beam path. When locking the IFO, one locks YARM first, then XARM. This means that it makes sense to look at YARM FF first rather than XARM.
3) XARM FF can't be done now since GUR2 is sketchy.
I'm planning on using this eLOG entry to document my Journey and Adventures (Chapter 2: YARM) to the OPTIMAL land of zero-seismic-noise (ZSN) at the 40m telescope.
Yesterday I adjusted the preweighting of my IIR fit to the transfer function of MC2, and also managed to reduce the number of poles and zeros from 8 to 6, giving a smoother rolloff. The bode plots are pictured here:
The predicted IIR subtraction was very close to the predicted FIR subtraction, so I thought these coefficients would lead to a better online filter.
However, the actual subtraction of the MCL was not as good and noise was injected into the Y arm.
The final comparison of the subtraction factors between the online and offline data showed that the preweighting, while it improved the offline subtraction, needs more work to improve the online subtraction also.
Since I will need to do transfer function measurements in order to implement FF for the arms and the MC2's yaw and pitch channels, I decided to practice this by replicating the transfer function measurement Eric did for MC2 to MCL. I followed his procedure and the data that I aquired for the TF looked as shown below,
About five minutes of data were taken (0.05 Hz resolution, 25 averages) by injecting noise from 1 to 100 Hz. The TF coherence looked as below,
Last night, I also worked on a "better" (an improvement, I thought) of the MISO Wiener filter (T240-X and T240-Y witnesses) IIR filter. The FF results are shown below:
Training data + Predicted FIR and IIR subtraction:
Online subtraction results:
Although the predicted FIR and IIR results are "better" than the previous MISO filter, the subtraction performance for this filter is marginally better if not worse (both peak at 15 dB, in slightly different regions).
Last night I performed some MISO FF on MCL using the T240-X and T240-Y as witnesses. Here are the results:
I've tweaked the ELOG code to allow uploading of PDFs by drag-and-drop into the main editor window. Once again we can bask in the glory of
In previous elogs, we saw that the X and Y spectra out of GUR2 (X end Guralp seismometer) looked strange (i.e. inconsistent with the GUR1 spectra).
This morning, Steve and I brought the handheld control unit to the Guralp to center the test mass, by adjusting the centering potentiometers inside the unit while monitoring the voltage readout corresponding to the DC mass position (manual has instructions).
At first glance, this seemed like the likely culprit, as the offsets for the horizontal directions were much larger than the vertical one. We zeroed all three to within a mV or two. Unfortuntately, the spectra look the same as they did 10 hours ago.
Since we already had the kit out, we checked the offsets for GUR1. Only the "East/West" had an offset over 50mV, so we zeroed that one, but left the others alone.
In my previous elog:11492, I stated that in order to improve the subtraction and reduce the injection of high frequency noise we want the filter's magnitude to have a 1/f rolloff.
I implemented this scheme on the filter SISO filter previously analyzed. The results are shown below.
The filters bode plot:
The nice 1/f rollof is the main change here. Everything else remained pretty much the same.
The predicted FIR and IIR subtractions:
Everything looks right but that hump at 8 Hz. I used 8 pairs of poles/zeros to get this subtraction.
The online MCL subtraction:
This looks better than I expected. One has to keep in mind that I ran this at 1 AM. I wonder how well this filter will do during the noisier hours of the day. The RMS at high frequencies doesn't look great, there will definitely be noise being injected into the YARM signal at high frequencies.
Measuring the YARM signal:
There is still noise being injected on YARM but it is definitely much better than the previous filter. I'm thinking about doing some IIR subtraction on the arms now to see if I can get rid of the noise that is being injected that way, but before I embark on that quest I will rething my prefiltering.
The plot below shows the ratio of the unfiltered versus filtered ASDs for the FIR and IIR subtraction predictions as well as for the measured online IIR subtraction. Positive dB means better subtraction.
Today I finished fitting the transfer function to a vectfit model for seismometers T240_X and T240_Y, and then used these to filter noise online from the mode cleaner.
The Bode plot for T240_X is in figure 1, and T240_Y is in figure 2. I made sure to weight the edges of the fit so that no DC coupling or excessive injection of high frequency noise occurs at the edges of the fit.
I used C1:IOO-MC_L_DQ as the first channel I filtered, with C1:IOO-MC_L_DQ(RMS) for RMS data. I took reference data first, without my filter on. I then turned the filter on and took data from the same channel again. The filtered data, plotted in red, subtracted from the reference and did not inject noise anywhere in the mode cleaner.
I also looked at C1:LSC-YARM_OUT_DQ and C1:LSC-YARM_OUT_DQ(RMS) for its RMS to see if noise was being injected into the Y-Arm when my filter was implemented. I took reference data here also, shown in blue, and compared it to data taken with the filter on. My filter, in pink, subtracted from the Y-Arm and injected no noise in the region up to 10 Hz, and only minimal noise at frequencies ~80 Hz. Frequencies this high are noisy and difficult to filter anyways, so the noise injection was minimal in the Y-Arm.
I’m working on a code to determine the Gaussianity of a PSD.
It can be difficult to distinguish between GW events and non-Gaussian noise, especially in burst searches. By characterizing noise Gaussianity, we can better recognize noise patterns and distinguish between GW events and noise.
What I did:
I analyzed an hour of S5 L1 data. First, I plotted a timeseries, just to see what I was working with. Then, I produced a PSD (technically, an ASD) for the timeseries using Welch’s method in Python.
I split the data segment into smaller time-chunks and then produced a PSD for each chunk. All PSDs were superimposed in one plot. Here’s a plot for 201 time-chunks of equal length:
For a specific frequency, I can view the spread in PSD value through the production of a histogram.
I’ve made histograms displaying varying PSD values for the 201 PSD plot at 100 Hz, 500 Hz, and 1kHz.
For Gaussian noise, an exponential decay plot is expected. I will continue this analysis by following the statistical method in Ando et al. 2003 to calculate specific values indicative of the Gaussianity of various distributions. I’ll then look at different periods of time in the S5 L1 data to find periods of time suggesting non-Gaussian behavior.
The wasp terminator came in today. He obliterated the known wasp nest.
We discovered a second wasp nest, right next to the previous one...
Jessica wasn't too happy the wasps weren't gone!
Last night we finally got some online subtraction going. The filter used is described in the post this eLOG is @eLOG 11488.
The results were as follow:
The filter worked as expected when subtracting noise out of MCL,
There is about a factor of 6 subtraction at the ~3Hz resonant peak. The static IIR filter predicted a factor of 6-7 subtraction of this peak as well.
The 1.2 Hz resenonant feature improved by a factor of 3. This should improve quite drastically when I implement the y-channel of the T240 seismo.
There is some high frequency noise being injected, not very noticeable, but present.
We then took a look at the power in the MC when the filter was on,
The power being transmitted in the cavity was not as stable as with the feedforward on. We believe that the filter is not at fault for this as Eric mentioned to me that the MC2 actuator lacked some sort of compensation that I need to understand a bit better.
YARM was then locked when the filter was on and we took a look at how it was doing. There was stationary sound arising from the locking of the YARM, leading us to believe that the filter might have injected some noise in the signal. IT DID.
The filter injected nasty high frequency noise at YARM from 11 Hz and on. This is to be expected since the filter did not roll off to zero at high frequencies. Implementing a 1/f rolloff should mitigate some of the injected noise.
Also, as one can see above, subtraction by around a factor of 2 or so, was induced by the mode cleaner feedforward subtraction.
After testing both the Conductive and Isolated front panels on the ALS delay line box using the actual beatbox and comparing this to the previous setup, I found that the conductive SMAs improved crosstalk the most. Also, as the old cables were 30m and the new ones are 50m, Eric gave me a conversion factor to apply to the new cables to normalize the comparison.
I used an amplitude of 1.41 Vpp and drove the following frequencies through each cable:
X: 30.019 MHz Y: 30.019203 MHz
which gave a difference of 203 Hz.
In the first figure, it can be seen that, for the old setup with the 30m cables, in both cables there is a spike at 203 Hz with an amplitude of above 4 m/s^2/sqrt(Hz). When the 50m cables were measured in the box with the conductive front panel, the amplitude drops at 203 Hz by a factor of around 3. I also compared the isolated front panel with the old setup, and found that the isolated front panel worse by a factor of just over 2 than the old setup. Therefore, I think that using the conductive front panel for the ALS Delay Line box will reduce noise and crosstalk between the cables the most.
Jessica will soon ELOG about some measurements suggesting that the conductive connector-ized ALS delay line enclosure is the way to go, when considering crosstalk between the delay lines. It is currently mounted and hooked up on the LSC rack, though I need to make a bunch of new SMA cables now that I think a semi-permanent arrangement has been reached.
I did a rough re-calibration of the phase tracker output, since the increased cable delay changes the degree/Hz gain. This was done by fitting a line to a slow sawtooth FM of the SRS DS345's (1Hz rate, 10kHz deviation, 30MHz carrier). This resulted in the following calibration updates
Again, this is a rough calibration. Nevertheless, it is not so surprising we don't get the 50m/30m = 4.4dB increase we would expect just from the lengths; the (I presume) increased cable loss matters. Also, the loss' frequency dependance is an additional reason that the phase tracker calibration is not constant over all frequencies.
I took spectra with the arms in IR lock, but didn't see any real improvement beyond a possible dip in the floor from 100-200Hz. This doesn't surprise me too much, however, since I don't believe that we are currently dominated by electronic noises that this gain increase would help overcome.
Last week, Koji mentioned the ALS phase noise added due to the post-cavity table motion the arm-transmitted green beams experience before hitting the beat PD. I should estimate the size of this effect for our situation.
Short and sweet of it:
I'm attaching a SISO IIR Wiener filter here for reference purposes that will go online either tonight or tomorrow evening. This is a first test to convince myself that I can get this to work, MISO IIR filters are close to being ready and will soon be employed.
This Wiener filter uses the STS-X channel as a witness and MCL as target. The bode plot for the filter is shown below,
The performance of the FIR and IIR Wiener filters and the ammount of subtraction achive for MCL is shown below,
Output from quack to be loaded with foton: filter.zip
The wasp nest will be removed tomorrow from from the out side of the east arm window.
The resonant frequency of the newly arrived gravity bee detector is not known.
Often when I come to manually lock the mode cleaner due to a long unlocked period, I find that the sliders are not in the state specified by the mcdown script. Furthermore, it's not the same channels every time; sometimes the servo gain is left high, sometimes the boosts are left on. I fear that some of the caput commands are failing to execute. Ugh.
This continues to happen. I believe the network latency boogeyman is to blame.
There was a long unlocked period because the enable switch for the MC servo fast path (FASTSW) was left off. Running the mcdown script fixed this, but included the error message:
Channel connect timed out: 'C1:IOO-MC_REFL_GAIN' not found.
CA Client Library: Ignored duplicate create channel response from CA server?
which means the IN1 gain didn't get touched. A second pass of the script produced no errors.
I'm thinking of adding some logic that if the autolocker has failed to lock for some period (5 minutes?), it should rerun mcdown.
In order to do online static IIR Wiener filtering one needs to do the following,
1) Get data for FIR Wiener filter witnesses and target.
2) Measure the transfer function (needs to be highly coherent, ~ 0.95 for all points) from the actuactor to the control signal of interest (ie. from MC2_OUT to MC_L).
3) Invert the actuator transfer function.
4) Use Vectfit or (LISO) to find a ZPK model for the actuator transfer and inverse transfer functions.
5) Prefilter your witness data with the actuator transfer function, to take into account the actuator to control transfer function when designing the offline Wiener FIR filter.
6) Calculate the frequency response for each witness from the FIR coefficients.
7) Vectfit the frequency reponses to a ZPK model, this is the FIR to IIR Wiener conversion step.
8) Now, either, divide the IIR transfer function by the actuator transfer function or more preferably, multiply by the inverse transfer function.
9) Use Quack to make SOS model of the IIR Wiener filter and inverse transfer function product that goes into foton for online implementation.
10) Load it into the system.
The block diagram below summarizes the steps above.
Koji had suggested that I sync up the two function generators to ensure that they have the same base frequency and so that crosstalk will actually appear at the expected frequency. After syncing up the two function generators, I drove the following frequencies through each cable:
X: 29.537 MHz Y: 29.5372 MHz
X: 29.545 MHz Y: 29.5452 MHz
Each time, the difference between the frequencies was 200 Hz, so if there was crosstalk, a spike should appear in the PSDs at 200 Hz when frequencies are being driven through both cables simulataneously, but not when just one is on. We very clearly see a spike at 200 Hz in both the X arm and the Y arm with the conductive SMAs, indicating crosstalk. For the front panel with isolated SMAs, we see a spike at 200 Hz when both frequencies are on, but it is much less pronounced than with the conductive SMAs. It seems as though there will be crosstalk using either panel, just less with the isolated SMAs.
Tonight, I've taken a bunch of data where the PRC is carrier locked and the ITM oplevs have the DC coupling FM turned on, as we use during locking. This is to inform new feedforward filters to stabilize the PRC angular motion, by using Wiener filtering with the POP QPD as the target, and local seismometers/accelerometers as witnesses. So far I've looked at the 1800 seconds leading up to GPS time 1122885600, but there has been plenty of locked time tonight if I need to retrieve more.
I've also measured the PRM ASC output torque -> POP QPD spot motion with high (>0.95) coherence from 0.1Hz to 10Hz.
Prefiltering so far consists of a 4th order elliptic LP at 5 Hz, with the target subtraction band being the 1-3Hz range.
With offline FIR filtering, the RMS pitch motion is reduced by a factor of 3 just with the STS1_X data. IIR fitting remains to be done.
The PRC yaw motion, which is marginally noisier, is a little more tangled up across X and Y.
Plots / filters forthcoming pending more analysis.
Since Chiara's onboard ethernet card has a reputation to be flaky in Linux, Koji suggested we could just buy a new ethernet card and throw it in there, since they're cheap.
I've installed a Intel EXPI9301CT ethernet card in Chiara, which detected it without problems. I changed over the network settings in /etc/networking/interfaces to use eth1 instead of eth0, restarted nfs and bind9, and everything looked fine.
Sadly, EPICS/network slowdowns are still happening. :(
I've fixed the ASC tab on the summary pages to populate the graphs with data without causing an error.
Motivation: The ASC tab was showing no data. It resulted in a name error when generated.
A name error indicates a bad channel name in the plot definition. I identified two errors in the code:
The plots are not processing without error. However, no titles or axis labels are present on the plots- I'll work on adding these.
Last night around 1AM, many of the the frontend models crashed due to an ADC timeout. (But none of the IOPs, and all the c1lsc models were fine.)
[1502036.695639] c1rfm: ADC TIMEOUT 0 46281 9 46153
[1502036.945259] c1pem: ADC TIMEOUT 0 56631 55 56695
[1502036.965969] c1mcs: ADC TIMEOUT 1 56706 2 56770
[1502036.965971] c1sus: ADC TIMEOUT 1 56706 2 56770
Then, simultaneously on c1ioo, c1iscex, and c1iscey. (Wed Aug 5 01:10:53 PDT 2015)
[1509007.391124] c1ioo: ADC TIMEOUT 0 46329 57 46201
[1509007.702792] c1als: ADC TIMEOUT 1 63128 24 63192
[2448096.252002] c1scx: ADC TIMEOUT 0 46293 21 46165
[2448096.258001] c1asx: ADC TIMEOUT 0 46669 13 46541
[1674945.583003] c1scy: ADC TIMEOUT 0 46297 25 46169
[1674945.685002] c1tst: ADC TIMEOUT 0 52993 1 52865
I'm still working on getting things back up and running. Just restarting models wasn't working, so I'm trying some soft reboots...
UPDATE: A soft reboot of all frontends seems to have worked,
Notes from tonight's work:
I've explored the beatnote fluctuations a bit further.
First, I realized that we already had a channel than functions much like an RF level monitor: the phase tracker Q output. I verified that indeed, the Q signal agrees with the RF monitor signals from the demod board within the phase tracker bandwidth. This simplifies things a little.
I also found that the Y beat suffers a fair bit less from these effects; which isn't too surprising given our experience with the alignment stability.
One possible caveat to my earlier conclusions is that the beatnote amplitude could be fluctuating due to real RIN of the green light transmitted through the cavity. In fact, this effect is indeed present, but can't explain all of the coherence. If it did, we would expect the DC green PDs (ALS-TR[X/Y]) to show the same coherence profile as the RF monitors, which they don't.
The next thing I was interested was whether the noise level predicted via coherence was realistic.
To this end, I implemented a least-squares subtraction of the RF level signal from the phase tracker output. I included a quadratic term of the RF power, but this turned out to be insiginficant.
Indeed, using the right gain, it is possible to subtract some noise, reproducing nearly the same spectrum as the coherence based estimate. The discrepency at 1Hz is possible from 1Hz cavity RIN, as suggested by the presence of some coherence with TRX.
However, this is actually kind of weird. In reality, I would've expected the coupling of RF level fluctuations to be more like a bilinear coupling; changing the gain of the mixer, rather than directly introducing a linearly added noise component. Maybe I just discovered the linear part, and the bilinear coupling is the left over low frequency noise... I need to think this over a little more.
Previously, I had gotten the same results for the conductive and the isolated front panels. Today, I sanded off the anodized part on the back of the conductive front panel. I checked afterwards with a mulitmeter to ensure that it was indeed conductive through all the SMA connectors.
I drove a frequency of 29.359 Hz through the X Arm cable and 29.3592 Hz through the Y Arm cable, giving a difference of 200 Hz. Previously, there would only be a spike in the Y Arm at the difference, while the X Arm did not change if the Y arm was on or off. Now that the panel is fully conductive, a spike can also be seen in the X arm, indicating that crosstalk may possibly be happening with this panel, now that the spike corresponds to both the X arm and Y arm. These results are only after one set of data. Tomorrow I'll take two more sets of data with this panel and do a more in depth comparison of these results to what had been previously seen.
I've added states to the summary pages to only show data for times at which one certain channel is above a specified threshold. So far, I've incorporated states for the IOO tab to show when the mode cleaner is locked.
You can see these changes implemented in the IOO tab of my personal summary pages for June 30: https://ldas-jobs.ligo.caltech.edu/~eve.chase/summary/day/20150630/ioo/.
I've written a description of how to add states to summary pages here: https://wiki-40m.ligo.caltech.edu/DailySummaryHelp#How_to_Define_and_Implement_States.
Atm1, New short-50" long cable was installed at ETMY end ( Y-station ) between Guralp-B ( MIT ) and granite base.
Interface box input 2 was left connected to cable 1 and input 1 to cable 2. This plot shows no change.
Atm2, Than I swapped the two long cables at the interface box
Now the signal seems to be ok <2 Hz,
>2 Hz some problem exist.
50" short cable
I will look for more bad soldering tomorrow. How many cables did she make?
We have to redo this cable also
Rana pointed out that another way to mitigate seismic motion at in the mode cleaner would be to look at the YAW and PITCH output channels of the WFS sensors that control the angular alignment of the mode cleaner.
I downloaded 45 mins of data from the following two channels:
And did some quick offline Wiener filtering with no preweighting, the results are shown in the PSD's below,
I'm quite surprised at the Wiener subtraction obtained for the YAW signal, it required no preweighting and there is about an order of magnitude improvement in our region of interest, 1-3 Hz. The PIT channel didn't do so bad either.
I tested both of the front panels (conductive and isolated SMAs) with the ALS Delay Line Box by driving extremely close frequencies through the cables. By doing this, we would expect that a spike would show up in the PSD if there was crosstalk between the cables.
In the plots below, for the conductive panel, the frequencies used were
X Arm: 22.329 MHz Y Arm: 22.3291 MHz
For the isolated panel, the frequencies were
X Arm: 22.294 MHz Y Arm: 22.2943 MHz
This gives a difference of 100 Hz for the conductive panel and 300 Hz for the isolated panel. Focusing on these areas of the PSD, it can be seen that in the Y Arm cable there is a very clear spike within 30 Hz of these differences when frequencies are being driven through both cables as opposed to the signal being in only the Y Arm. In the X Arms, the noise in general is higher when both cables are on, but there is no distinct spike at the expected frequencies. This indicates that some sort of crosstalk is probably happening due to the strong spikes in the Y Arm cables.
However, I wonder how much of the low frequency noise can be explained by instability of the beat alignement on the PSL table, and how this might be quantified.
I followed my hunch, and the truth comes out.
I recalled that the aLIGO demod board has a handy DB9 output on the back panel for the detected power at the RF and LO inputs. I hooked this up into the BEATY ADC channels while checking the ALSX spectrum in IR lock.
This is assuredly the limiting factor in our ALS sensitivity.
Note: I'm calling the fluctuations of the beatnote amplitude "RF Amplitude RIN," to put things in reasonble units. I haven't looked up the board's conversion of dBm to V, but the LO should be around 0dBm in this measurement.
The coherence between the phase tracker output and the LO amplitude is significant over a broad range, mostly dipping where real cavity motion peeks up into the spectrum.
Also, the feature from 10-100Hz in the RIN spectrum is one I've often seen directly in ALS spectra when beatnote alignement is bad or the beatnote frequency is high, convincing me further that this is what's to blame.
So: what do we do? Is there anything we can do to make the green alignment more stable?
ALS is not currently limited by the demod board or whitening electronics.
The noise budget in the green locking paper shows the main noise sources to be these two, plus the residual fluctuations of the green PDH loop.
So, one next step is AUX PDH noise budget.
Yesterday, I put together a few measurements to asses whether the new demod board has moved us in the right direction. Specifically I measured the output of the phase tracker in the following states, adjusting the phase tracker gain to maintain a ~2kHz UGH (but no boost on):
Results: The beat frequency spectrum is above the measured demod board and whitening chassis/ADC noise at all frequencies. It's a little close at 10Hz.
One nice feature is that the beat spectra are far more similar to each other than they used to be. RMS noise is in the 300-400Hz range, which isn't mindblowing, but not terrible. On the order of 50 pm for each arm. Most of this comes from below 10Hz.
Another thing to note is that, when we switch in the 50m cables, we should win a fair bit of Hz/V gain and push down these noises futher. (We're currently using 30m cables.)
By looking at some coherences, we can attribute some of the noise when IR locked to both colors of PDH loops.
Specifically, the coherence with the Green PDH error implicates the residual frequency noise of the AUX laser above a few hundred Hz, whereas the feature from 20-50Hz is probably real cavity motion, not ALS sensing noise. Some of the 1-3Hz noise is from real suspension/stack resonances too.
If it turns out that we do want to push the demod board noise down further, we could think about increasing the RF amplification. Driving the board harder translates directly to better noise performance. The 60Hz harmonics aren't so exciting, but not the end of the world.
Data files are attached, if you're in to that sort of thing.
I've switches the ALS, ASC, and LSC plots on the summary pages from plotting raw frames, to plotting minute trends, instead. Now, the plots contain information, instead of being completely blank, but data is not recorded on the plots after 12UTC.
Typically, I make changes to the summary pages on my own version of the pages, found at https://ldas-jobs.ligo.caltech.edu/~eve.chase/summary/day/, where I change the summary pages for June 30 and then import such changes into the main summary pages.
Please check the spectra. If something is wrong, please swap the cables between X and Y in order to see if the cable is still the issue. I believe the cable was nicely made as I carefully checked the connection twice or more during and after the soldering work.
Today it was synced at 5AM but that was all.
controls@nodus|minute > pwd
controls@nodus|minute > ls -l 11222|tail
-rw-r--r-- 1 controls controls 35521183 Jul 29 21:59 C-M-1122264000-3600.gwf
-rw-r--r-- 1 controls controls 35509281 Jul 29 22:59 C-M-1122267600-3600.gwf
-rw-r--r-- 1 controls controls 35511705 Jul 29 23:59 C-M-1122271200-3600.gwf
-rw-r--r-- 1 controls controls 35809690 Jul 30 00:59 C-M-1122274800-3600.gwf
-rw-r--r-- 1 controls controls 35752082 Jul 30 01:59 C-M-1122278400-3600.gwf
-rw-r--r-- 1 controls controls 35927246 Jul 30 02:59 C-M-1122282000-3600.gwf
-rw-r--r-- 1 controls controls 35775843 Jul 30 03:59 C-M-1122285600-3600.gwf
-rw-r--r-- 1 controls controls 35648583 Jul 30 04:59 C-M-1122289200-3600.gwf
-rw-r--r-- 1 controls controls 35643898 Jul 30 05:59 C-M-1122292800-3600.gwf
-rw-r--r-- 1 controls controls 35704049 Jul 30 06:59 C-M-1122296400-3600.gwf
controls@nodus|minute > ls -l 11223|tail
-rw-r--r-- 1 controls controls 35696854 Jul 30 08:02 C-M-1122300000-3600.gwf
-rw-r--r-- 1 controls controls 35675136 Jul 30 08:59 C-M-1122303600-3600.gwf
-rw-r--r-- 1 controls controls 35701754 Jul 30 09:59 C-M-1122307200-3600.gwf
-rw-r--r-- 1 controls controls 35718038 Jul 30 10:59 C-M-1122310800-3600.gwf
LDAS Minute trend:
[koji.arai@ldas-pcdev3 C-M-11]$ pwd
[koji.arai@ldas-pcdev3 C-M-11]$ ls -l |tail
-rw-r--r-- 1 1001 1001 35518238 Jul 29 19:59 C-M-1122256800-3600.gwf
-rw-r--r-- 1 1001 1001 35514930 Jul 29 20:59 C-M-1122260400-3600.gwf
-rw-r--r-- 1 1001 1001 35521183 Jul 29 21:59 C-M-1122264000-3600.gwf
-rw-r--r-- 1 1001 1001 35509281 Jul 29 22:59 C-M-1122267600-3600.gwf
-rw-r--r-- 1 1001 1001 35511705 Jul 29 23:59 C-M-1122271200-3600.gwf
-rw-r--r-- 1 1001 1001 35809690 Jul 30 00:59 C-M-1122274800-3600.gwf
-rw-r--r-- 1 1001 1001 35752082 Jul 30 01:59 C-M-1122278400-3600.gwf
-rw-r--r-- 1 1001 1001 35927246 Jul 30 02:59 C-M-1122282000-3600.gwf
-rw-r--r-- 1 1001 1001 35775843 Jul 30 03:59 C-M-1122285600-3600.gwf
-rw-r--r-- 1 1001 1001 35648583 Jul 30 04:59 C-M-1122289200-3600.gwf
Koji soldered new 50" long cable for the Y station.
The refreshed ALS didn't look so bad today (elog forthcoming), so I decided to give PRFPMI locking a shot tonight. I was able to hold the PRMI while swinging through resonsance, but transitions to RF signals failed. Demod angles / whitening gains/ etc. etc. all need to be rechecked
Some little things here and there that got cleaned up...
Eric downloaded MC2 to MCL transfer function data (H) as well as its inverse, MCL to MC2 (Hinv). He also downloaded new MCL and MC2 data.
I used vectfit to fit the MC2 to MCL transfer function,
The ZPK parameters for this fit were,
Zeros 1278.36719876674 + 0.00000000000000i
-100.753249679343 + 0.00000000000000i
-18.6014192997845 + 13.0294910760217i
-18.6014192997845 - 13.0294910760217i
Poles -1.11035771175328 + 7.03549674098987i
-1.11035771175328 - 7.03549674098987i
-18.8655320274072 + 0.00000000000000i
-690.294337433234 + 0.00000000000000i
Using the above vectfit model, I filtered the raw MC2 signal to get 'MCL'. The PSD's of the raw MCL data and the filtered MC2 result is shown below,
The lack of accuracy of the transfer function at replicating MCL at frequencies lower than 0.7Hz is expected, the vectfit model I generated fails to follow accurately the raw transfer function data. My question: Does it matter? My guess: Probably not. In order to mitigate seismic noise from the mode cleaner we are mainly concerened with the 1-3 Hz region.
I also used vectfit to fit the transfer function for MCL to MC2,
This one was harder to fit accurately for some reason, I could do it with four pairs of zeros and poles but it took some preweighting.
The ZPK parameters for the above fit were,
Zeros 0.173068278283995 + 0.00000000000000i
0.995140531040529 + 0.0268079821980457i
0.995140531040529 - 0.0268079821980457i
0.894476816129099 + 0.00000000000000i
Poles -19.9566906920707 + 18.0649464375308i
-19.9566906920707 - 18.0649464375308i
-109.275971483008 + 0.00000000000000i
-1791.88947801703 + 0.00000000000000i
Similarly, using this ZPK model, I filtered the MCL signal to get 'MC2'. I plotted the PSD for the MC2 signal and the filtered MCL to get,
Again, the lack of accuracy of the filtered MC2 at replicating MCL below 0.7 Hz and above 12 Hz is due to the inverse transfer function failing to converge in these ranges.